# Achieving Pareto Optimality in Games via Single-bit Feedback

### v1

@@ 1 Introduction
Achieving socially optimal coordination in multi-agent systems is challenging under severe communication limitations. For example, in wireless sensor networks or dynamic spectrum access, agents optimize a global objective (a social welfare) using only local observations and minimal feedback. Prior research shows that even payoff-based learning rules can guide agents toward efficient outcomes in generic $n$-player games without inter-agent communication Marden et al. (2012); Pradelski and Young (2012). However, these approaches typically lack effective finite-time performance guarantees. This gap motivates the need for a new learning mechanism that is both communication-efficient and backed by rigorous convergence guarantees.
@@
We propose Single-Bit Coordination Dynamics for Pareto-Efficient Outcomes (SBC-PE)—a multi-agent learning algorithm requiring only one bit per agent per round. The method applies to arbitrary finite games and maximizes a weighted sum of local payoffs. SBC-PE follows an explore-then-commit paradigm: agents first explore by randomly sampling actions, then commit to those most often aligned with collective satisfaction. At each step, agents broadcast a single-bit signal generated probabilistically from their realized utilities and parameters $(w_{i},\lambda_{i},\epsilon)$; this minimal feedback is the only exchanged information. Despite such limited communication—far lower than typical schemes—the dynamics provably drives the system to the exact socially optimal outcome. This low burden makes SBC-PE especially attractive for domains with bandwidth constraints or intermittent connectivity.
@@
Communication efficiency is a central challenge in distributed learning, with strategies such as event-triggering, sparsification, and quantization proposed to mitigate bandwidth and energy costs Cao et al. (2023). Quantization-based methods are especially appealing, with single-bit protocols representing the most communication-efficient extreme. Prior work shows that compressive and reduced-dimension diffusion strategies achieve near full-information performance while lowering communication Sayin and Kozat (2014, 2013). Similarly, consensus and optimization methods using only the sign of relative state differences demonstrate that one-bit messages suffice for distributed agreement and optimization Zhang et al. (2019); Chen et al. (2011). Despite such compression, single-bit strategies maintain comparable convergence rates and sometimes attain exact solutions Zhang et al. (2019, 2018). Building on this line, our contribution shows that in arbitrary finite games, a single satisfaction bit per agent suffices to preserve communication efficiency while guaranteeing exact convergence to the socially optimal outcome with finite-time regret bounds.
@@
SBC-PE provably converges to the exact optimum in a decentralized way. For sufficiently large horizon $T$, the expected total regret is logarithmic, $O(\log T)$, with a precise bound in Theorem 4. This result and our numerical simulations show that SBC-PE identifies the welfare-maximizing joint action $a^{*}$ efficiently in finite time.
@@
Earlier distributed learning approaches aimed for efficiency but did not quantify performance. For example, the Game of Thrones algorithm achieves decentralized coordination without communication, but is analyzed only in terms of asymptotic regret and limited to collusion-based resource allocation Bistritz and Leshem (2021). Notably, the payoff-based rules can be stochastically stable at Pareto-efficient solutions Marden et al. (2012) or optimal equilibria Pradelski and Young (2012). However, they do not provide finite-time rate guarantees on optimality. By contrast, SBC-PE combines minimal signaling with provable finite-time guarantees.
@@ 2 Problem Formulation
We consider a multi-agent game with agents $N=\{1,2,\dots,n\}$. Each agent $i$ selects an action $a_{i}\in A_{i}$, where $A_{i}$ is its finite action set. The joint action is $a=(a_{1},\dots,a_{n})\in A=\prod_{i=1}^{n}A_{i}$, over a finite horizon $T$. Each agent $i$ then receives a utility $u_{i}(a)$. To ensure fairness, each agent has a minimum utility threshold $\lambda_{i}$, requiring
$u_{i}(a)\geq\lambda_{i},\quad\forall i\in N$ (1)
so that no agent is entirely neglected when welfare is maximized.
The system objective is to maximize weighted sum:
$\max_{a\in A}W(a)=\sum_{i=1}^{n}w_{i}u_{i}(a),$ (2)
where $w_{i}>0$ denotes the weight of agent $i$. This aligns with distributed welfare optimization in learning-in-games and cooperative control.
Strategy updates rely only on (i) agent’s realized utility and (ii) one-bit feedback from others. This enforces a decentralized, payoff-based learning rule where agents cannot observe actions or utilities of others, yet achieve efficient communication. Agents communicate only via one-bit messages.
@@ 3 Algorithm
Initialize: exploration horizon $K$, content action counter $c_{i}(a_{i})=0\ \forall a_{i}\in A_{i}$, weight $w_{i}$, threshold $\lambda_{i}$, parameter $\epsilon\in(0,1)$
@@
   choose action $a_{i}(t)$ uniformly from $A_{i}$
   observe own realized utility $u_{i}(t)=u_{i}(a(t))$
   compute message
$m_{i}(t)=\begin{cases}\mathbb{I}\{u_{i}(t)>\lambda_{i}\}&\text{w.p. }\epsilon^{\,1-w_{i}u_{i}(t)},\\ 0&\text{otherwise}.\end{cases}$ (3)
   broadcast $m_{i}(t)$ to all other agents
   if $m_{i}(t)=1\ \forall i\in N$ then
     $c_{i}(a_{i}(t))=c_{i}(a_{i}(t))+1$
@@
for each time step $t>K$ do
   play the component $a_{i}^{*}$ corresponding to agent $i$
@@
We present a decentralized learning algorithm to maximize the social welfare objective under the communication constraint in (1). Algorithm 1 has two phases: an exploration phase of length $K$, followed by exploitation.
Our approach draws on Pradelski and Young (2012) and the payoff-algorithm in Bistritz and Leshem (2021), which achieve efficient outcomes without inter-agent communication (the latter using a small internal state). In contrast, our model is deliberately simple: it requires no state values and relies only on a one-bit satisfaction signal, yielding an easily implementable mechanism with $\epsilon$-bounds and convergence to Pareto optimality.
The Single-Bit Coordination Dynamics for Pareto-Efficient Outcomes (SBC-PE) is an explore-then-commit procedure (Algorithm 1). In the exploration phase, agents play random actions for $K$ rounds, observe utilities, and broadcast a one-bit signal $m_{i}(t)$ as in (3). Each agent keeps counters $c_{i}(a_{i})$, incrementing when all broadcast $m_{i}(t)=1$ generated stochastically from $u_{i}$, $w_{i}$, $\lambda_{i}$, and $\epsilon\in(0,1)$. After exploration, each selects $a_{i}^{*}=\operatorname*{arg\,max}_{a_{i}\in A_{i}}c_{i}(a_{i})$. In the exploitation phase, agents repeatedly play $a_{i}^{*}$, so the joint action $a^{*}=(a_{1}^{*},\dots,a_{n}^{*})$ converges to a Pareto-efficient outcome. Each agent acts independently, using only local utility and one-bit feedback.
@@ 4 Results
In this section, we analyze the regret bounds of the Algorithm 1, and specify the necessary conditions for which this holds.
Define $A_{\lambda}\coloneqq\{\,a\in A:u_{i}(a)>\lambda_{i}\ \forall i\in N\,\}$ as the feasible set of joint actions. For notational simplicity, we restrict attention to $A_{\lambda}$ in the proof. Let $a^{*}\in\operatorname*{arg\,max}_{a\in A_{\lambda}}W(a)$, $\Delta_{1}=W(a^{*})-\max_{a\neq a^{*}}W(a)$, $\Delta=\max_{a}(W(a^{*})-W(a))$, and $M\coloneqq|A_{\lambda}|$.
We need the following assumptions for the regret bound: Assumption 1.
@@
    The social welfare maximizer $a^{*}$ is unique,
@@
    $\epsilon<(M+\delta)^{-1/\Delta_{1}}$ for some $\delta>0$.
Let $c_{t}(a)\coloneqq\sum_{k=0}^{t}\mathbb{I}_{\{a(k)=a\}}\mathbb{I}_{\{m_{i}(t)=1,\ \forall i\in N\}}$ denote the count of each joint action $a\in A$ up to time $t$ for which all agents are content. Similarly, let $c_{i,t}(a_{i})\coloneqq\sum_{k=0}^{t}\mathbb{I}_{\{a_{i}(k)=a_{i}\}}\mathbb{I}_{\{m_{i}(t)=1,\ \forall i\in N\}}$ denote the count of individual actions $a_{i}\in A_{i}$ for agent $i\in N$ under contentment. If $\epsilon$ is sufficiently small, the most frequent joint action and the individual agents’ most frequent actions coincide.
For each $a\in A_{\lambda}$, define a Bernoulli random variable
$\rho(a)=\begin{cases}1&\text{if }m_{i}(t)=1,\ \forall i\in N,\\ 0&otherwise\end{cases}$
Then $P(\rho(a)=1)=\prod\epsilon^{1-w_{i}u_{i}(a)}=\epsilon^{,n-W(a)}$. By the Strong Law of Large Numbers (SLLN),
$\lim_{t\rightarrow\infty}\tfrac{c_{t}(a)}{t}=P(\rho(a)=1)=\epsilon^{,n-W(a)}\quad\textit{a.s.}$
Since $\epsilon<1$,
$\operatorname*{arg\,max}_{a\in A_{\lambda}}\epsilon^{n-W(a)}=\operatorname*{arg\,max}_{a\in A_{\lambda}}W(a)=a^{,}$
where the last equality follows from the uniqueness of the welfare maximizer. Define $\theta\in\mathbb{R}^{M}$ with entries $\theta(a)=P(\rho(a)=1)$. For each agent $i$, the marginalization of $\theta$ is
$\theta_{i}(\tilde{a}_{i})\coloneqq\sum_{\{a\in A_{\lambda}:\ a_{i}=\tilde{a}_{i}\}}\theta(a).$ (4)
Then, by the SLLN, we have $\lim_{t\rightarrow\infty}c_{i,t}(a_{i})/t=\theta_{i}(a_{i})$. Remark 1. Normalizing $\theta$ yields a softmax distribution over the social welfare values of the joint actions:
$\bar{\theta}(a)\coloneq\frac{\theta(a)}{\|\theta\|_{1}}=\frac{(1/\epsilon)^{W(a)}}{\sum_{\tilde{a}\in A_{\lambda}}(1/\epsilon)^{W(\tilde{a})}}$.
The next lemma shows how well the optimal joint action $a^{*}$ is separated from all others. Lemma 1. Under Assumption 4, we have
$\theta(a^{*})-\xi>\sum_{a\neq a^{*}}(\theta(a)+\xi),$ (5)
where $\xi\coloneqq\frac{\delta}{M}\epsilon^{N-\max_{\tilde{a}\neq a^{*}}W(a)}$
@@
$\theta(a^{*})-M\xi=\epsilon^{N-W(a^{*})}-\delta\epsilon^{N-\max_{\tilde{a}\neq a^{*}}W(a)}=\epsilon^{N-\max_{\tilde{a}\neq a^{*}}W(a)}(\epsilon^{-\Delta_{1}}-\delta)$ $>M\epsilon^{N-\max_{\tilde{a}\neq a^{*}}W(a)}$
$>\sum_{a\neq a^{*}}(\theta(a)).$
@@
During exploration, agents must sample each joint action multiple times to estimate $\theta$. Let $\widehat{\theta}_{t}(a)\coloneqq c_{t}(a)/t$ and $\widehat{\theta}_{i,t}(a_{i})\coloneqq c_{i,t}(a_{i})/t$ be estimates of $\theta$ and $\theta_{i}$. In practice, agents do not observe $\widehat{\theta}_{t}$; agent $i$ only sees $\widehat{\theta}_{i,t}$.
If each entry of $\theta$ can be estimated within error $\xi$, the following proposition shows that agents can identify the action required to achieve the optimal social welfare.
Proposition 1. If Assumption 4 holds, and $|\widehat{\theta}_{t}(a)-\theta(a)|<\xi$ for all $a\in A_{\lambda}$, then we have
$\{\operatorname*{arg\,max}_{a_{i}\in A_{i}}\widehat{\theta}_{i,t}(a_{i})\}_{i=1}^{n}$ $=\operatorname*{arg\,max}_{a\in A_{\lambda}}\theta(a)=a^{*}.$ (6)
Proof: From the definition of $\widehat{\theta}_{i,t}$, and $\widehat{\theta}_{t}$, we have
$\widehat{\theta}_{i,t}(a_{i}^{*})=\sum_{\{a\in A_{\lambda}:\ a_{i}=a_{i}^{*}\}}\widehat{\theta}_{t}(a)>\widehat{\theta}_{t}(a^{*})>\theta(a^{*})-\xi.$
Using Lemma 4, we have
$\theta(a^{*})-\xi>\sum_{a\neq a^{*}}(\theta(a)+\xi)>\sum_{a\neq a^{*}}\widehat{\theta}_{t}(a)>\sum_{\{a\in A_{\lambda}:\ a_{i}\neq a_{i}^{*}\}}\widehat{\theta}_{t}(a)>\widehat{\theta}_{i,t}(\tilde{a}_{i}),\ \forall\tilde{a}_{i}\neq a_{i}^{*},$
Hence, we can say that $\operatorname*{arg\,max}_{a_{i}\in A_{i}}\widehat{\theta}_{i,t}(a_{i})=a_{i}^{*}$, and we obtain (6). $\square$
@@
Remark 3. If the joint action is observable to all agents, the dynamics still guarantee coordination without Assumptions 1-(ii) and 1-(iii), converging to one of the social maximizers.
Theorem 1 (Main Result) For any multi-agent game, with a finite horizon length T, if Assumption 4 holds, for large enough T, the expected total regret is bounded by
$\bar{R}<\tfrac{M\Delta}{\xi^{2}}\left(1+\left(1-\tfrac{M}{T\xi^{2}}\right)\log\left(\tfrac{2TM\xi^{2}}{M-1}\right)\right)=O(log(T)).$
Proof: With exploration length $K$, let $l_{K}(a)$ denote the number of times joint action $a$ is played. This is a binomial random variable. By Chernoff bound Chernoff (1952), for any $a\in A_{\lambda}$,
$P\left(l_{K}(a)<\tfrac{K}{2M}\right)<e^{-\frac{K}{8M}}.$ (7)
Using the Hoeffding bound, the estimation error of $\widehat{\theta}(a)$ for any $a\in A_{\lambda}$ satisfies
$P\left(\widehat{\theta}_{K}(a)-\theta(a)\geq\xi,\Big|,l_{K}(a)>\tfrac{K}{2M}\right)\leq 2e^{-\frac{K}{M}\xi^{2}}.$ (8)
Then, by (7), (8), and union bound,
$\sum_{a\in A_{\lambda}}P(\widehat{\theta}_{K}(a)-\theta(a)\geq\xi)<\tfrac{2Me^{-\frac{K}{M}\xi^{2}}}{1-Me^{-\frac{K}{8M}}}.$ (9)
Assume $K>16M\ln(M)$ to simplify the denominator in (9). This is feasible for large $T$ since $M$ is constant. Then
$P\big(\widehat{\theta}_{K}(a)-\theta(a)\leq\xi,\ \forall a\in A_{\lambda}\big)>1-\beta,$ (10)
$\text{with}\quad\beta\coloneqq\psi e^{-\frac{K}{M}\xi^{2}},\quad\psi\coloneqq\tfrac{2M^{2}}{M-1}.$
By Proposition 4 and (10), agents play the best joint action in the exploitation phase with probability $1-\beta$ and a worse joint action with probability $\beta$. Thus, the expected regret is:
$\bar{R}_{K}<K\Delta+\beta(T-K)\Delta.$
If we find the optimal $K^{*}$ that minimizes $K\Delta+\lambda T\Delta$, by equating the derivative to zero, we have
$K^{*}=\tfrac{M}{\xi^{2}}\log{\left(\tfrac{T\psi\xi^{2}}{M}\right)}.$
Choosing $K=K^{*}$, $\beta=\frac{M}{T\xi^{2}}$, we have
$\bar{R}<\tfrac{M\Delta}{\xi^{2}}\left(1+\left(1-\tfrac{M}{T\xi^{2}}\right)\log\left(\tfrac{2TM\xi^{2}}{M-1}\right)\right),$
and, we can bound the expected regret. $\square$

### v2

@@ Abstract
Keywords: Pareto efficiency, single-bit communication, regret analysis, multi-agent system.
@@ 1 Introduction
Achieving socially optimal coordination in multi-agent systems is challenging under severe communication limitations. For example, in wireless sensor networks or dynamic spectrum access, agents optimize a global objective (a social welfare) using only local observations and minimal feedback. Prior research shows that even payoff-based learning rules can guide agents toward efficient outcomes in generic $n$-player games without inter-agent communication (Marden et al., 2012; Pradelski and Young, 2012). However, these approaches typically lack effective finite-time performance guarantees. This gap motivates the need for a new learning mechanism that is both communication-efficient and backed by rigorous convergence guarantees.
@@
We propose Single-Bit Coordination Dynamics for Pareto-Efficient Outcomes (SBC-PE)—a multi-agent learning algorithm requiring only one bit per agent per round. The method applies to arbitrary finite games and maximizes a weighted sum of local payoffs. SBC-PE follows an explore-then-commit paradigm: agents first explore by randomly sampling actions, then commit to those most often aligned with collective satisfaction. At each step, agents broadcast a single-bit signal generated probabilistically from their realized utilities and parameters; this minimal feedback is the only exchanged information. Despite such limited communication—far lower than typical schemes—the dynamics provably drives the system to the exact socially optimal outcome. This low burden makes SBC-PE especially attractive for domains with bandwidth constraints or intermittent connectivity.
@@
Communication efficiency is a central challenge in distributed learning, with strategies such as event-triggering, sparsification, and quantization proposed to mitigate bandwidth and energy costs (Cao et al., 2023). Quantization-based methods are especially appealing, with single-bit protocols representing the most communication-efficient extreme. Prior work shows that compressive and reduced-dimension diffusion strategies achieve near full-information performance while lowering communication (Sayin and Kozat, 2014, 2013). Similarly, consensus and optimization methods using only the sign of relative state differences demonstrate that one-bit messages suffice for distributed agreement and optimization (Zhang et al., 2019; Chen et al., 2011). Despite such compression, single-bit strategies maintain comparable convergence rates and sometimes attain exact solutions (Zhang et al., 2019, 2018). Building on this line, our contribution shows that in arbitrary finite games, a single satisfaction bit per agent suffices to preserve communication efficiency while guaranteeing exact convergence to the socially optimal outcome with finite-time regret bounds.
@@
SBC-PE provably converges to the exact optimum in a decentralized way under mild assumptions. For sufficiently large horizon $T$, the expected total regret is logarithmic, $O(\log T)$, with a precise bound in Theorem 4. This result and our numerical simulations show that SBC-PE identifies the welfare-maximizing joint action efficiently in finite time.
@@
Earlier distributed learning approaches aimed for efficiency but did not explicitly quantify the performance. For example, the Game of Thrones algorithm achieves decentralized coordination without communication, but is analyzed only in terms of asymptotic regret and limited to collusion-based resource allocation (Bistritz and Leshem, 2021). Notably, the payoff-based rules can be stochastically stable at Pareto-efficient solutions (Marden et al., 2012) or optimal equilibria (Pradelski and Young, 2012). However, they do not provide finite-time rate guarantees on optimality. By contrast, SBC-PE combines minimal signaling with provable finite-time guarantees.
@@ 2 Problem Formulation
We consider a multi-agent game with agents $N=\{1,\dots,n\}$. Each agent $i$ selects an action $a_{i}\in A_{i}$, where $A_{i}$ is its finite action set. The joint action is $a=(a_{1},\dots,a_{n})\in A=\prod_{i=1}^{n}A_{i}$. Each agent $i$ then receives a utility $u_{i}(a)$. Agents collectively aim to maximize the social welfare:
$W(a)=\sum_{i=1}^{n}w_{i}\cdot u_{i}(a),$ (1)
where $w_{i}>0$ is the weight of agent $i$. Each agent can have local constraints such as tolerable utility threshold $\lambda_{i}$, requiring
$u_{i}(a)>\lambda_{i},\quad\forall i\in N.$ (2)
Agents interact over $T$ stages. Across these stages, their strategy updates rely only on (i) agent’s realized local utility and (ii) one-bit feedback from others. This enforces a decentralized, payoff-based learning rule where agents cannot observe actions or utilities of others, yet achieve efficient outcomes with minimal communication.
@@ 3 Algorithm
Require: exploration length $K$, local weight $w_{i}>0$, local threshold $\lambda_{i}$, common parameter $\epsilon\in(0,1)$
Initialize: local counter $c_{i}(a_{i})=0\ \forall a_{i}\in A_{i}$
@@
   choose action $a_{i}^{t}\sim\textrm{Uniform}(A_{i})$
   observe local utility $u_{i}^{t}=u_{i}(a^{t})$
   construct message
$m_{i}^{t}=\begin{cases}\mathbb{I}_{\{u_{i}^{t}>\lambda_{i}\}}&\text{with probability }\epsilon^{1-w_{i}\cdot u_{i}^{t}},\\ 0&\text{otherwise}.\end{cases}$ (3)
   broadcast $m_{i}^{t}\in\{0,1\}$ to all other agents
   if $m_{j}^{t}=1\ \forall j\in N$ then
     increment $c_{i}(a_{i}^{t})=c_{i}(a_{i}^{t})+1$
@@
for each time step $t=K+1,\ldots,T$ do
   play $\bar{a}_{i}$
@@
We present a decentralized learning algorithm to maximize the social welfare objective under the feasibility constraint (2). Algorithm 1 has two phases: an exploration phase of length $K$, followed by exploitation for TKT-K stages.
Our approach draws on (Marden et al., 2012; Pradelski and Young, 2012; Bistritz and Leshem, 2021). They use internal states and aspiration levels while requiring sufficiently small design parameters for performance guarantees without explicit bounds. In contrast, our model is deliberately simple: it requires no state values or aspiration levels and relies only on a one-bit satisfaction signal, yielding an easily implementable mechanism with performance guarantees for Pareto optimality under explicit bounds on the design parameter $\epsilon$.
The Single-Bit Coordination Dynamics for Pareto-Efficient Outcomes (SBC-PE) is an explore-then-commit procedure (Algorithm 1). In the exploration phase, agents play random actions for $K$ rounds, observe utilities, and broadcast a one-bit signal $m_{i}^{t}$ stochastically generated as in (3). Each agent $i$ keeps counters $c_{i}(a_{i})$, incrementing when all broadcast $\{m_{j}^{t}=1\}_{j\in N}$. After exploration, agent $i$ selects $\bar{a}_{i}\in\operatorname*{arg\,max}_{a_{i}\in A_{i}}c_{i}(a_{i})$. In the exploitation phase, agent $i$ repeatedly plays the selected action $\bar{a}_{i}$, so the joint action $\bar{a}=(\bar{a}_{1},\dots,\bar{a}_{n})$ achieves the Pareto-efficient outcome. We highlight that each agent acts independently, using only local utility and one-bit feedback.
@@ 4 Results
In this section, we focus on regret-based analysis to evaluate the performance of Algorithm 1. Given the joint actions $a^{t}$ played at stages $t=1,\ldots,T$, the cumulative regret relative to the optimal social welfare is defined by
$R_{T}=T\max_{a\in A_{\lambda}}\{W(a)\}-\sum_{t=1}^{T}W(a^{t}),$ (4)
where $A_{\lambda}\coloneqq\{\,a\in A:u_{i}(a)>\lambda_{i}\ \forall i\in N\,\}$ is the feasible set of joint actions.
We make the following assumptions for the regret bound:
Assumption 1.
@@
    There exists a unique social welfare maximizer:
    $a^{*}=\operatorname*{arg\,max}_{a\in A_{\lambda}}\{W(a)\}.$ (5)
@@
    For some $\delta>0$, we have
    $\epsilon<(M+\delta)^{-1/\Delta_{1}},$ (6)
    where $M\coloneqq|A_{\lambda}|$ and
    $\Delta_{1}\coloneqq\max_{a\in A_{\lambda}}\{W(a)\}-\max_{\begin{subarray}{c}a\in A_{\lambda}:\\ a\neq a^{*}\end{subarray}}\{W(a)\}.$ (7)
Define $m^{t}\coloneqq\prod_{j\in N}m_{j}^{t}$, taking value $1$ when each signals $m_{j}^{t}=1$. Then, by (1) and (3), the independent randomization of messages yields
$P(m^{t}=1\mid a^{t}=a)=\left\{\begin{array}[]{ll}\epsilon^{n-W(a)}&\mbox{if }a\in A_{\lambda}\\ 0&\mbox{otherwise}\end{array}\right.$. (10)
Correspondingly, at each stage $t$, agents jointly play $a$ and signal content with probability
$\theta(a)\coloneqq P(a^{t}=a,m^{t}=1)=\left\{\begin{array}[]{ll}\epsilon^{n-W(a)}/|A|&\mbox{if }a\in A_{\lambda}\\ 0&\mbox{otherwise}\end{array}\right$.
due to the uniform exploration.
Remark 1. The content-endorsed actions are distributed by
$P(a^{t}=a\mid m^{t}=1)=\left\{\begin{array}[]{ll}\frac{e^{-W(a)}}{\sum_{\tilde{a}\in A_{\lambda}}e^{-W(\tilde{a})}}&\mbox{if }a\in A_{\lambda}\\ 0&\mbox{otherwise}\end{array}\right$. (13)
Since $\epsilon<1$, Assumptions 1-(i) and 1-(ii) yield
$\operatorname*{arg\,max}_{a\in A}\{\theta(a)\}=\operatorname*{arg\,max}_{a\in A_{\lambda}}\{W(a)\}=a^{*}.$ (14)
The following lemma shows how well $\theta(a^{*})$ is separated from all others $\theta(a)$ for $a\neq a^{*}$.
Lemma 1. Under Assumption 4, we have
$\theta(a^{*})-\xi>\sum_{\begin{subarray}{c}a\in A_{\lambda}:\\ a\neq a^{*}\end{subarray}}(\theta(a)+\xi),$ (15)
where
$\xi\coloneqq\frac{\delta}{|A|M}\cdot\epsilon^{n-\max_{\tilde{a}\neq a^{*}}W(a)}.$ (16)
@@
$\theta(a^{*})-M\xi$ $=\frac{1}{|A|}\epsilon^{n-\max_{\tilde{a}\neq a^{*}}W(a)}(\epsilon^{-\Delta_{1}}-\delta)$ (17)
$>\frac{M}{|A|}\epsilon^{n-\max_{\tilde{a}\neq a^{*}}W(a)}$ (18)
$>\sum_{\begin{subarray}{c}a\in A_{\lambda}:\\ a\neq a^{*}\end{subarray}}\theta(a).$ (19)
@@
As an unbiased estimate of $\theta(a)$, we introduce
$\widehat{\theta}^{t}(a)\coloneqq\frac{1}{t}\sum_{k=1}^{t}\mathbb{I}_{\{a^{k}=a,m^{k}=1\}}.$ (20)
By the Strong Law of Large Numbers, we have $\widehat{\theta}^{t}(a)\rightarrow\theta(a)$ as $t\rightarrow\infty$ almost surely for each $a\in A$. Furthermore, the marginalization yields
$\theta_{i}(\tilde{a}_{i})\coloneqq\sum_{\begin{subarray}{c}a\in A:\\ a_{i}=\tilde{a}_{i}\end{subarray}}\theta(a)\quad\mbox{and}\quad\widehat{\theta}_{i}^{t}(\tilde{a}_{i})\coloneqq\sum_{\begin{subarray}{c}a\in A:\\ a_{i}=\tilde{a}_{i}\end{subarray}}\widehat{\theta}^{t}(a).$ (21)
If each entry of $\theta$ can be estimated within error $\xi$, the following proposition shows that agents can locally identify the action required to achieve the optimal social welfare.
Proposition 1. For a given stage $t$, if Assumption 4 holds, and $|\widehat{\theta}^{t}(a)-\theta(a)|<\xi$ for all $a\in A$, then we have
$\left(\operatorname*{arg\,max}_{a_{i}\in A_{i}}\left\{\widehat{\theta}_{i}^{t}(a_{i})\right\}\right)_{i=1}^{n}$ $=\operatorname*{arg\,max}_{a\in A}\{\theta(a)\}=a^{*}.$ (22)
Proof: From the definition of $\widehat{\theta}_{i}^{t}$, and $\widehat{\theta}^{t}$, we have
$\widehat{\theta}_{i}^{t}(a_{i}^{*})=\sum_{\begin{subarray}{c}a\in A:\\ \ a_{i}=a_{i}^{*}\end{subarray}}\widehat{\theta}^{t}(a)>\widehat{\theta}^{t}(a^{*})>\theta(a^{*})-\xi.$ (23)
On the other hand, Lemma 4 yields
$\theta(a^{*})-\xi>\sum_{a\neq a^{*}}(\theta(a)+\xi)$ $>\sum_{a\neq a^{*}}\widehat{\theta}^{t}(a)>\widehat{\theta}_{i}^{t}(\tilde{a}_{i})$ (24)
for all $\tilde{a}_{i}\neq a_{i}^{*}$. Hence, $\widehat{\theta}_{i}^{t}(a_{i}^{*})>\widehat{\theta}_{i}^{t}(\tilde{a}_{i})$ for all $\tilde{a}_{i}\neq a_{i}^{*}$, and we obtain (22). $\square$
@@
Remark 3. If the joint action is observable to all agents, the dynamics still guarantee convergence to one of the social maximizers without Assumptions 1-(ii) and 1-(iii).
The following theorem provides $O(\log(T))$ bound on the expected regret (4), where the expectation is taken with respect to the randomness induced by the uniform exploration and stochastic feedback.
Theorem 1 (Main Result) For any $n$-agent game, if Assumption 4 holds and the exploration length is $K=\tfrac{\log(4MT\xi^{2})}{2\xi^{2}}$, then the expected total regret is bounded by
$\mathrm{E}[R_{T}]<\frac{\Delta}{2\xi^{2}}\left(1+\log(4MT\xi^{2})\left(1-\frac{1}{2T\xi^{2}}\right)\right),$ (25)
where $\Delta\coloneqq\max_{a}(W(a^{*})-W(a))$, $M=|A_{\lambda}|$, and $\xi$ is as described in (16).
Proof: Based on the randomness on $\mathbb{I}_{\{a^{t}=a,m^{t}=1\}}\in\{0,1\}$ and (20), the Hoeffding bound yields that
$P\left(\left|\widehat{\theta}^{K}(a)-\theta(a)\right|\geq\xi\right)\leq 2e^{-2K\xi^{2}}$ (26)
for each $a$. Then, using the union bound, we obtain
$P\left(\left|\widehat{\theta}^{K}(a)-\theta(a)\right|\geq\xi\mbox{ for some }a\in A\right)\leq 2Me^{-2K\xi^{2}}.$
Correspondingly, we have
$P\left(\left|\widehat{\theta}^{K}(a)-\theta(a)\right|\leq\xi,\ \forall a\in A\right)\geq 1-\beta,$ (27)
where $\beta\coloneqq 2Me^{-2K\xi^{2}}$. By Proposition 4 and (27), agents play the best joint action $a^{*}$ in the exploitation phase with probability at least $1-\beta$, and a worse joint action $a\neq a^{*}$ with probability at most $\beta$. Thus, the expected regret is:
$\mathrm{E}[R_{T}]\leq K\Delta+\beta(T-K)\Delta<K\Delta+\beta T\Delta.$
The optimal $K^{*}$ minimizing $K\Delta+\beta T\Delta$ is given by
$K^{*}=\tfrac{1}{2\xi^{2}}\log(4MT\xi^{2}).$
Choosing $K=K^{*}$, we obtain $\beta=\frac{1}{2T\xi^{2}}$ and (25), which completes the proof. $\square$